# 6G: The Intelligent Network of Everything – A Comprehensive Vision, Survey, and Tutorial

### v2

@@ VIII-J Security Technologies for 6G
VIII-J1 Network Security Architecture
@@
    New types of security and privacy threats were introduced for 5G due to its expanded capabilities and a broad range of novel applications. In this respect, 3GPP’s Release 15 introduced an advanced security architecture, with features such as access-agnostic authentication, enhanced subscription privacy, user plane integrity protection, network slice-specific authentication/authorization, and advanced authentication/authorization between network functions [638]. During the entire evolution of 5G, its security architecture will be constantly updated, considering newly added technology features and extended capabilities. Currently, a security and privacy threat paradigm is studied for 6G in academia and industry, while paying attention to the emerging 6G technologies, services, and applications.
@@ IX Defining Features for 6G
XI 7G Vision: A High-Level Sketch
In this section, we place 6G into a wider perspective by discussing the post-6G era and sketching the first high-level vision of 7G. As each generation, 6G will eventually reach its limits and the next generation (7G) will gradually take over. Most likely, this evolution will follow the same ten year cycle seen in the past. While 6G is expected to dominate in the 2030s, we envision that the 2040s will be the 7G era. The sooner we start visioning the next generation, the more prepared we are and the more time we have to develop it. Although 6G is in the development process and many years from the deployment phase, we can already see some major trends that evolve beyond 6G. While wireless will evolve toward unlimited, AI will expand vertically (deeper) and horizontally (broader). In addition, the range of connected devices will become wider and wider. Based on these trends, we can envision the next major disruption after 6G.
What comes to the disruptions provided by mobile networks, it seems that the major disruptions occur in the cycles of two generations, i.e., 1G/2G, 3G/4G, and 5G/6G. In this cycle, the first generation introduces the new disruption, whereas the second one makes it boom. For example, 1G introduced mobile telephony, while 2G freed its potential. Mobile internet was first enabled by 3G and became mainstream in the 4G era. Currently, 5G is introducing mobile and intelligence and it is expected that 6G will take the mobile intelligence to the next level. This evolution will most likely continue after 6G. We envisage that the next major disruption in the post-6G era will be the 7G/8G-enabled mobile hyperverse. Mobile hyperverse will rely on the cornerstones of 7G. We identify these three fundamental pillars as unlimited wireless, deep intelligence (DI), and the Internet of Intelligence (IoI). These elements are vital to enable 7G-level applications. We identify some possible applications in three categories, i.e., human-DI interactions, deeply intelligent world, and interconnected intelligent ecosystems. In the following, we provide further details on the aforementioned disruption, elements, and applications.
TABLE XIV: A speculative evolution from 6G to 7G
@@
We envision that the mobile hyperverse will be the next evolution phase after mobile intelligence, creating real-time intelligent, interconnected, interactive, and immersive entities that can ultimately integrate the digital, physical, biological, and even abstract worlds. The scale of integration can vary from nano to global range. Roughly speaking, the hyperverse can be seen as an extension of the metaverse. Mobile hyperverse can take many forms and serve different purposes. Ultimately, this evolution might lead to the Internet of Hyperverses. At a high level, mobile hyperverses can be divided into three categories: public, non-public, and hybrid. Public hyperverses are intended for consumers, whereas the non-public ones are for industrial, organizational, and enterprise purposes. Hybrid hyperverses are a mix of both, being partly public and partly non-public. From a human perspective, mobile hyperverse will enrich human interactions with other people, devices, systems, and applications, revolutionizing the human experience of the world. We envisage that the mobile hyperverse will become mainstream when all seamlessly interconnected parts of the entire ecosystem have reached a certain level of maturity, including networks, devices, user interfaces, applications, and intelligence. Mobile hyperverse requires ubiquitous, super-quality, and intelligent wireless network infrastructure; smart, lightweight, and affordable XR device entities; smooth and easy to use super-precision user interfaces; fully immersive applications exploiting all human senses; and advanced AI deeply integrated into all parts of the ecosystem. The 7G-enabled mobile hyperverse will create countless application and business opportunities at all levels of the future society.
@@
Unlimited wireless refers to virtually unlimited wireless resources in diverse dimensions, including connectivity, computation, control, information, positioning, sensing, and energy. As AI will expand vertically and horizontally in the future, we refer to this advanced level of AI as DI. Due to DI, IoE will evolve toward IoI, which can be seen as a network of deeply intelligent entities. What comes to the 7G-level applications, human-DI interactions refer to diverse types of interactions between people and deeply intelligent entities. Examples include an augmented human, brain-DI interfaces, personal DI, telepathic communications, and holographic worlds. A deeply intelligent world refers to a world enhanced by DI at all levels and scales. Example applications in this domain include global intelligent systems, global monitoring/management/prediction entities, and a deeply intelligent society. Interconnected intelligent ecosystems refer to large-scale entities on the ground, in the air, and at the sea, exploiting wireless, DI, and IoI technologies. The evolution from 6G to 7G is summarized in Table XIV.
XII Conclusion

### v3

@@ IX-J SECURITY TECHNOLOGIES FOR 6G
IX-J1 NETWORK SECURITY ARCHITECTURE
@@
    New types of security and privacy threats were introduced for 5G due to its expanded capabilities and a broad range of novel applications. In this respect, 3GPP Release 15 introduced an advanced security architecture, with features such as access-agnostic authentication, enhanced subscription privacy, user plane integrity protection, network slice-specific authentication/authorization, and advanced authentication/authorization between network functions [647]. During the entire evolution of 5G, its security architecture will be constantly updated, considering newly added technology features and extended capabilities. Currently, a security and privacy threat paradigm is studied for 6G in academia and industry, while paying attention to the emerging 6G technologies, services, and applications.
@@ X DEFINING FEATURES FOR 6G
XII 7G VISION: A HIGH-LEVEL SKETCH
In this section, we place 6G into a wider perspective by discussing the post-6G era and sketching the first high-level vision of 7G. As each generation, 6G will eventually reach its limits and the next generation (7G) will gradually take over. Most likely, this evolution will follow the same ten year cycle seen in the past. While 6G is expected to dominate in the 2030s, we envision that the 2040s will be the 7G era. Since 7G will be significantly more complicated than 6G, it will take more time to develop it. The sooner we begin visioning 7G, the more prepared we are and the more time we have to develop it. Although 6G is in the development process and many years from the deployment phase, we can already see some major trends that evolve beyond 6G. While wireless will evolve toward unlimited, AI will expand vertically (deeper) and horizontally (broader). In addition, the range of connected devices will become wider and wider. Based on these trends, we can envision the next major disruption after 6G.
What comes to the disruptions provided by mobile networks, it seems that the major disruptions occur in the cycles of two generations, i.e., 1G/2G, 3G/4G, and 5G/6G. In this cycle, the first generation introduces the new disruption, whereas the second one makes it boom. For example, 1G introduced mobile telephony, while 2G freed its potential. Mobile internet was first enabled by 3G and became mainstream in the 4G era. Currently, 5G is introducing mobile and intelligence and it is expected that 6G will take mobile intelligence to the next level. This evolution will most likely continue after 6G. We envisage that the next major disruption in the post-6G era will be the 7G/8G-enabled mobile hyperverse. Mobile hyperverse will rely on the cornerstones of 7G. We identify these three fundamental pillars as unlimited wireless, deep intelligence (DI), and the Internet of Intelligence (IoI). These elements are vital to enable 7G-level applications. We identify some possible applications in three categories, i.e., human-DI interactions, deeply intelligent world, and interconnected intelligent ecosystems. In the following, we provide further details on the aforementioned disruption, elements, and applications.
TABLE XVI: A speculative evolution from 6G to 7G
@@
We envision that mobile hyperverse will be the next evolution phase after mobile intelligence, creating real-time intelligent, interconnected, interactive, and immersive entities that can ultimately integrate the digital, physical, biological, and even abstract worlds. The scale of integration can vary from nano to global range. Early signs of this kind of evolution can already be seen in Nokia’s 6G vision, which aims to connect the digital, physical, and human worlds [49]. Roughly speaking, the hyperverse can be seen as an extension of the metaverse. Mobile hyperverse can take many forms and serve different purposes. Ultimately, this evolution might lead to the Internet of Hyperverses. At a high level, mobile hyperverses can be divided into three categories: public, non-public, and hybrid. Public hyperverses are intended for consumers, whereas the non-public ones are for industrial, organizational, and enterprise purposes. Hybrid hyperverses are a mix of both, being partly public and partly non-public. From a human perspective, mobile hyperverse will enrich human interactions with other people, devices, systems, and applications, revolutionizing the human experience of the world. We envisage that the mobile hyperverse will become mainstream when all seamlessly interconnected parts of the entire ecosystem have reached a certain level of maturity, including networks, devices, user interfaces, applications, and intelligence. Mobile hyperverse requires ubiquitous, super-quality, and intelligent wireless network infrastructure; smart, lightweight, and affordable XR device entities; smooth and easy to use super-precision user interfaces; fully immersive applications exploiting all human senses; and advanced AI deeply integrated into all parts of the ecosystem. The 7G-enabled mobile hyperverse will create countless application and business opportunities at all levels of the future society.
@@
Unlimited wireless refers to virtually unlimited wireless resources in diverse dimensions, including connectivity, information, positioning, sensing, and energy. As AI will expand vertically and horizontally in the future, we refer to this advanced level of AI as DI. Due to DI, IoE will evolve toward IoI, which can be seen as a network of deeply intelligent entities. What comes to the 7G-level applications, human-DI interactions refer to diverse types of interactions between people and deeply intelligent entities. Examples include an augmented human, brain-DI interfaces, personal DI, telepathic communications, and holographic worlds. A deeply intelligent world refers to a world enhanced by DI at all levels and scales. Example applications in this domain include global intelligent systems, global monitoring/management/prediction entities, and a deeply intelligent society. Interconnected intelligent ecosystems refer to large-scale entities on the ground, in the air, and at the sea, exploiting wireless, DI, and IoI technologies. Overall, our 7G vision can be summarized in a few words: ”Deeply Intelligent Network of Hyperverses”. The evolution from 6G to 7G is summarized in Table XVI.
XIII CONCLUSION